OpenStudy (anonymous):

What is n ^-1/7 over n^2/7×n^-5/7

OpenStudy (jdoe0001):

\(\Large \bf \cfrac{n^{-\frac{1}{7}}}{n^{\frac{2}{7}}\cdot n^{-\frac{5}{7}}}?\)